The function f : x ı→ a + b cos x, is defined for 0 ≤ x ≤ 2π. Given that f(0) = 10 and that = 1 , find

(i) the values of a and b, (ii) the range of f, (iii) the exact value of
1
Expert's answer
2020-06-16T16:35:07-0400

"f(x)=a+b\cos x\\
f(0)=a+b\cos 0=10\to max\\
a+b=10\\
f(\pi)=a+b\cos \pi=1\to min\\
a-b=1\\
\left\{\begin{matrix}
 a +b=10 \\
 a-b=1
\end{matrix}\right.\\
i)\\
a=5.5\\
b=4.5\\
ii)\\
\cos x\in[-1,1]\\
4.5\cos x\in[-4.5,4.5]\\
f(x)=5.5+4.5\cos x\in[1,10]\\
iii)\\
f(x)=5.5+4.5\cos x\\
f(\frac{\pi}{2})=5.5+4.5\cdot\cos\frac{\pi}{2}=5.5\\
f(\frac{\pi}{3})=5.5+4.5\cdot\cos\frac{\pi}{3}=7.75"